OpenStudy (linaminas):

4(x+5) - 9 = 11 - 2(x -3)

OpenStudy (mariamikayla):

4(x+5)-9=11-2(x-3) 4x+20-9=11-2x+6 4x+11= 11 -2x +6 4x= -2x+6 4x+2x=6 6x=6 x=6/6=1

OpenStudy (linaminas):

You're a genius !

OpenStudy (linaminas):

So did you just subtract 11 from both side's?

OpenStudy (linaminas):

In the 4th step? @mariamikalya

OpenStudy (mariamikayla):

yes :))

OpenStudy (linaminas):

But how does the -2 become positive?

OpenStudy (mariamikayla):

because if we move a therm from the other part of the equal it changes it's sign

OpenStudy (linaminas):

Oh OK thank you